a ball was pushed by a force of 17.27 newton going to east. and travels a certain distance of 0.05678 kilometer in 0.033 minute. find the ball's momentum


1
Expert's answer
2022-05-27T09:38:57-0400
"d=v_0t+at^2/2,\quad v_0=0"

"a=2d/t^2=2*(56.78\rm\: m)/(1.98\:\rm s)^2=28.97\:\rm m/s^2"

"v_f=v_0+at=0+28.97*1.98=57.35\:\rm m/s"

"m=\frac{F}{a}=\frac{17.27\:\rm N}{28.97\:\rm m/s^2}=0.5961\:\rm kg"

The final momentum of the ball

"p_f=mv_f\\
=0.5961\:\rm kg\times 57.35\: m/s=34.19\: kg\cdot m/s"